tawneeharrison:

Jane Smart buys a new SUV. The price, including tax, is $22,500.00. She finances the vehicle over 60 months after making a $2,000 down payment. The true annual interest rate is 12%. What are Jane's monthly payments (principal plus interest)?

Vocaloid:

@sillybilly123 when you get a chance can you look at this

sillybilly123:

The amount of credit here is \(22,500 - 2,000 = 20,500 ~ USD\) The Monthly Instalment to pay this down is therefore: \(MI = P\cdot {\dfrac {r(1+r)^{n}}{(1+r)^{n}-1}}\) With \(P = 20,500\), \( r = \frac{12pc}{12} = 0.01\) and \(n = 5 \times 12 = 60\) \(=20,500 \cdot {\dfrac {0.01(1+0.01)^{60}}{(1+0.01)^{60}-1}} = USD 456.01\)